OpenStudy (anonymous):

solve this equation please: x^3+27=0 a medal will be given!

OpenStudy (anonymous):

x^3=-27 x=-3

OpenStudy (anonymous):

x^3 = -27 x = -3

OpenStudy (anonymous):

\[x ^{3}+27 = 0\] \[x ^{3}=-27\] \[\sqrt[3]{x}=-3\]

OpenStudy (anonymous):

u rock!!

OpenStudy (phi):

there is no easy way to get a cube root, unless you are allowed to use a calculator. Without a calculator, it is a good idea to memorize a few of the smaller cubes: 1 1 2 8 3 27 4 64 5 125 10 1000

OpenStudy (phi):

these questions are to separate out people who go that extra mile...

OpenStudy (anonymous):

Lol.